OpenStudy (anonymous):

If the midpoint between (x, 6) and (-9, 14) is (8, 10), find the value of x.

OpenStudy (anonymous):

the midpoint between two points (x1,y1) and (x2,y2) is

OpenStudy (anonymous):

\[(\frac{x_1+ x_2}{2},\frac{y_1+ y_2}{2})\]

OpenStudy (anonymous):

If the midpoint is at 8, and the end points are at x and -9, then (x +(-9))/2 = 8 x - 9 = 16 x = 25. For the second, if L is the midpoint, then seqments CL and LD are congruent. 1/3x + 8 = 2/3x +(-4) 12 = 1/3 x x = 36 So each segment is 20 and CD is 40. source: http://answers.yahoo.com/question/index?qid=20101101091658AApT457

OpenStudy (anonymous):

Thanks again @Eric_Playsinacoolway

OpenStudy (anonymous):

np!